Suppose that there are only two goods produced in the economy, Call center services and banking

services. Prices (P), quantities (Q) and the number of workers (W) occupied in the production of

each good for year 1 and year 2 are given by the following table :

Year 1 Year 2

Two

products

P1 Q1 W1 P2 Q2 W2

Call centre 10 100 50 12 100 50

Banking 10 200 50 12 230 60

a. What is nominal GDP in each year? ( 2 marks)

b. Using year 1 prices, what is real GDP in year 2? What is the growth rate of real GDP?

( 2 marks)

c. What is the rate of inflation using the GDP deflator? ( 2 marks)

d. Using year 1 prices, what is real GDP per worker in year 1 and year 2? What is labour

productivity growth between year 1 and year 2 for the whole economy?




1
Expert's answer
2021-12-02T10:37:41-0500

Solution:

a.). Nominal GDP is derived by multiplying the current year's quantity output by the current market price.

Call center services:

Year 1 = 10 "\times" 100 = 1,000

Banking Services:

Year 1 = 10 "\times" 200 = 2,000

Nominal GDP Year 1 = 1,000 + 2,000 = 3,000

 

Call center services:

Year 2 = 12 x 100 = 1,200

Banking Services:

Year 2 = 12 "\times" 230 = 2,760

Nominal GDP Year 1 = 1,200 + 2,760 = 3,960

 

b.). Real GDP is derived by using base-year prices and multiplying them by current year quantities for all the goods and services produced in an economy.

Call center services:

Year 2 = 10 "\times" 100 = 1,000

Banking Services:

Year 2 = 10 "\times" 230 = 2,300

Real GDP Year 2 = 1,000 + 2,300 = 3,300

 

The growth rate of Real GDP = (Real GDP in Year 2 – Real GDP in Year 1)/ Real GDP in Year 1

 

Growth rate of Real GDP = (3,300 – 3,000)/3000 = 300/3000 = 10%

Growth rate of Real GDP = 10%

 

Labor productivity growth Year 1 = Total output/Number of workers = 3,000/100 = 60

Labor productivity growth Year 2 = Total output/Number of workers = 3,960/110 = 30